On June 1, Parson Assoc. sold equipment to Arleo and agreed to accept a 3-month, $62,000,...

  1. On June 1, Parson Assoc. sold equipment to Arleo and agreed to accept a 3-month, $62,000, 10% interest-bearing note in payment at a time when the prevailing rate of interest for similar transactions was 10%. When the note was collected upon maturity, Parson would recognize interest revenue of:
Answer #1

Answer.

Interest Revenue recognized at the time of maturity = Face Value of Note \times Interest Rate \times Time

= $62,000 \times 0.10 \times 3/12 = $1,550

Thus, Parson would recognize interest revenue of $1,550 at the time maturity of Note. This interest revenue is presented in income statement under the head other income (non operating Income).